The perimeter of a rectangular field is 360 yards. If the width of the field is 84 yards, what is its length?
Answer by checkley77(12844)

P=2L+2W
360=2L+2*34
360=2L+68
2L=360-68
2L=292
L=292/2
L=146 YDS. IS THE LENGTH.
PROOF:
360=2*146+2*34
360=292+68
360=360